# Parallel Batch Scheduling With Incompatible Job Families Via Constraint Programming

### v2

@@ VIII Concluding remarks
This paper addressed the incompatible case of parallel batch scheduling, where compatible jobs belong to the same family, and jobs of different families cannot be processed together in the same batch. In the context of the diffusion area of the SM process, batches cannot be interrupted mid-process and jobs must be synchronized within their batch. The NS model by Ham et al. [18] for this problem does not synchronize the jobs within their batch, allowing interruptions mid-process. To remedy this limitation, the paper presented three new CP models, including two adaptions of existing models and a novel RS model. The AS model of Nascimento et al. [8] was extended to handle multiple incompatible job families and satisfy the batch size requirements. The S model of Cakici et al. [10] was also extended to the incompatible case. Finally, the paper proposed a new RS model that exclusively uses global constraints and adds redundancy to the problem structure. The computational experiments demonstrate that the proposed RS model dominates all the other models: it finds solutions to all the instances, proves optimality in the largest amount of instances, produces solutions that are significantly better than other models, especially in the largest instances and achieves the best gaps in the least amount of time. Current research includes using CP for serial batching, and future research includes extending the efficient arc-flow formulations for compatible p-batch [16, 29, 28, 30] to the incompatible case and embedding the proposed formulations in larger scheduling contexts [40, 41]
@@ Acknowledgements
The authors would like to express their gratitude to Amira Hijazi for her support and helpful discussions.

### v3

@@ VIII Concluding remarks
This paper addressed the incompatible case of parallel batch scheduling, where compatible jobs belong to the same family, and jobs of different families cannot be processed together in the same batch. The state-of-the-art CP model for this problem—the Aligned (A) model—relies on state functions, global constraints, and alignment parameters available only in a commercial CP solver, making it solver-dependent. To overcome this limitation, we have presented four new CP models that can be implemented in other (open-source) solvers:
  • The Automaton (AU) model, which relies on automaton constraints, available in various open-source solvers.
  • The Assign-and-Schedule (AS) model, which extends the model by Nascimento et al. [8] to handle multiple incompatible job families and includes additional constraints to enforce batch size requirements. This model distinctly separates assignment decisions (handled by binary variables) from batch sequencing decisions (handled by CP scheduling logic).
  • The Synchronized (S) model, which extends the model by Cakici et al. [10] to accommodate incompatible families by replacing binary assignment variables with interval variables that are synchronized with batch sequencing decisions.
  • The Redundant Synchronized (RS) model, which exclusively uses global constraints for both assignment and scheduling decisions and introduces redundancy into the problem structure.
Extensive computational experiments on 1,000 instances under two objectives (TWCT and $C_{\max}$) and across two solvers (CP Optimizer and CP-SAT) showcase the implementation flexibility of our proposed models. Our results reveal that:
  • The A model consistently produces high-quality solutions (0% gap) but is limited by its reliance on proprietary solver features only available in the commercial solver CP Optimizer.
  • The clear separation between binary assignment and batch scheduling decisions enables the AS model to excel in the open-source solver CP-SAT. Thanks to the satisfiability components of the solver, which efficient handle binary structures, the AS model achieves very small gaps and even proves optimality in isolated cases.
  • The same separation that benefits the AS model in CP-SAT, however, undermines its performance in CP Optimizer, where global constraints lead to stronger propagation.
  • The AU model shows promising scalability for moderate-sized instances, while the RS model, with its redundant structure, offers a competitive trade-off between propagation strength and computational efficiency.
  • Incorporating symmetry-breaking (SB) constraints in the AS model under CP Optimizer is beneficial for proving optimality when time permits, but under tight time limits, these constraints become harder to satisfy and negatively impact solution quality.
  • In CP-SAT, the SB constraints in the AS model, implemented via reification, introduce overhead in larger instances, thereby reducing the quality of the solutions obtained.
  • The SB constraints in the RS model negatively impact its performance because they add additional overhead that complicates the propagation process, leading to more frequent search failures and increased solution times in larger instances.
  • Minimizing the makespan is considerably easier than minimizing the TWCT, mainly because the makespan objective is insensitive to the processing order and job weights. Instead, it focuses solely on the maximum completion time across all jobs, simplifying the scheduling decisions and often allowing models to prove optimality within the time limit.
By delivering additional solver-independent CP formulations, this paper not only extends the state-of-the-art for incompatible p-batch scheduling but also opens avenues for integrating these models into broader manufacturing scheduling contexts. Current research includes using CP for serial batching, while future work will focus on extending efficient arc-flow formulations for compatible p-batch [16, 31, 30, 32] to the incompatible case and embedding the proposed formulations within larger scheduling frameworks [47, 48].
@@ Acknowledgements
The authors would like to express their gratitude to Amira Hijazi for her support and helpful discussions. Additionally, to Petr Vilím, Eray Cakici, and the anonymous reviewers for their helpful comments that greatly improved the quality of the paper.